# A Survey of 3D Reconstruction with Event Cameras

### v2

@@
Abstract
Event cameras are rapidly emerging as powerful vision sensors for 3D reconstruction, uniquely capable of asynchronously capturing per-pixel brightness changes. Compared to traditional frame-based cameras, event cameras produce sparse yet temporally dense data streams, enabling robust and accurate 3D reconstruction even under challenging conditions such as high-speed motion, low illumination, and extreme dynamic range scenarios. These capabilities offer substantial promise for transformative applications across various fields, including autonomous driving, robotics, aerial navigation, and immersive virtual reality. In this survey, we present the first comprehensive review exclusively dedicated to event-based 3D reconstruction. Existing approaches are systematically categorised based on input modality into stereo, monocular, and multimodal systems, and further classified according to reconstruction methodologies, including geometry-based techniques, deep learning approaches, and neural rendering techniques such as Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS). Within each category, methods are chronologically organised to highlight the evolution of key concepts and advancements. Furthermore, we provide a detailed summary of publicly available datasets specifically suited to event-based reconstruction tasks. Finally, we discuss significant open challenges in dataset availability, standardised evaluation, effective representation, and dynamic scene reconstruction, outlining insightful directions for future research. This survey aims to serve as an essential reference and provides a clear and motivating roadmap toward advancing the state of the art in event-driven 3D reconstruction.
Index Terms:
3D Reconstruction, Event Camera, Neuromorphic Vision, Event-based Vision, Neural Radiance Fields, 3D Gaussian Splatting
footnotetext: Corresponding author. Email: vincent.qu@sydney.edu.au
I Introduction
Figure 1: Roadmap of 3D reconstruction with event cameras. It shows the development from event-based geometry to neural 3D rendering.
3D reconstruction is a fundamental technique that transforms 2D observations of real-world scenes, objects, or simulated environments into accurate 3D models [1]. Typical outputs from these reconstruction processes include depth maps [2], point clouds [3], meshes [4], and voxel representations [5]. Numerous sensing technologies are used to capture the essential data required for 3D reconstruction, including conventional RGB cameras [6], RGB-D cameras [7], structured light sensors [8], and LiDAR systems [9, 10]. However, each of these traditional sensors has inherent limitations: conventional RGB cameras struggle under extreme lighting conditions and rapid motion scenarios, whereas active sensors such as LiDAR and structured-light systems are typically bulky and require substantial physical space [10]. Consequently, event cameras have emerged as an attractive alternative or complementary sensing modality, addressing many of these drawbacks.
Event cameras, also referred to as neuromorphic cameras, silicon retinas, or dynamic vision sensors, are bio-inspired sensors designed to asynchronously capture brightness changes rather than producing images at a fixed frame rate [11]. Unlike conventional RGB cameras, each pixel in an event camera independently detects and reports intensity changes, effectively acting as an individual sensor element [12]. When the brightness at a pixel exceeds a predefined threshold, the event camera generates event data consisting of the pixel’s coordinates, timestamp, and the polarity of the brightness change. Due to their unique sensing principle, event cameras exhibit outstanding advantages, such as high temporal resolution, low latency, and robustness to motion blur and challenging lighting conditions [11]. These properties have driven extensive research on event cameras across various vision tasks, including object detection [13], recognition [14], segmentation [15], and tracking [16]; video enhancement tasks such as super-resolution [17], frame interpolation [18], and deblurring [19]; as well as complex spatiotemporal modeling tasks including simultaneous localisation and mapping (SLAM) [20], visual odometry (VO) [21], and notably, 3D reconstruction. Event-based vision technologies have demonstrated considerable potential in diverse applications, ranging from autonomous driving [22], robotics [23], and unmanned aerial vehicles (UAVs) [24] to industrial monitoring [25], etc.
@@ I Introduction
Leveraging their distinct sensing capabilities, event cameras have gained significant attention in recent years for performing 3D reconstruction. It is a domain traditionally dominated by standard cameras since the 1990s [26]. Event-based 3D reconstruction approaches now span from occupancy mapping [27] and real-time dynamic scene reconstruction [28], to achieving high-fidelity rendering of complex scenes [29]. Figure 1 illustrates the roadmap and evolution of event-based 3D reconstruction techniques. Initial efforts utilising event cameras for 3D reconstruction began emerging prominently in the 2010s [30, 31]. Existing reconstruction strategies employ stereo and monocular event camera setups [32, 33], alongside multimodal approaches integrating event data with other sensory modalities [34]. Recent advances further integrate sophisticated neural rendering methods, such as Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS), to enable unprecedented levels of realism [35, 29].
Figure 3: Publication timeline in categories (from 2015 to May 2025)
Figure 4: Categorisation of methods in the survey
@@
Notably, a complete pipeline from raw event data acquisition to final 3D model reconstruction should be able to output a representable 3D structure, either in the form of a directly modellable 3D representation or at least a depth map that can be readily converted into a 3D point cloud (as in some depth estimation methods [36, 37]). However, numerous earlier works that lack a final 3D output, such as those in SLAM [20], VO [21], and 3D perception [38], can be considered precursor steps toward 3D reconstruction, but should not be regarded as complete event-to-3D model reconstruction pipelines. In Figure 2, publications dedicated to event-based 3D reconstruction have significantly increased since 2015, experiencing rapid growth particularly in the last three years, underscoring the burgeoning interest and research value in this area. Since the choice of sensing configuration heavily influences the subsequent reconstruction pipeline, existing methods are typically categorised according to the camera setup: stereo, monocular, and multimodal approaches.
@@
Despite the steadily growing number of research publications, there currently exists no dedicated survey for event-driven 3D reconstruction. Three comprehensive but broader surveys have briefly touched upon aspects of 3D reconstruction using event cameras. Gallego et al. (2020) provided a wide-ranging overview of event cameras [26], which included a brief section on 3D reconstruction. However, rapid advances have rendered parts of that review outdated. More recently, Chakravarthi et al. (2024) categorised event camera research tasks [39], offering only limited insights into 3D reconstruction. Similarly, Zheng et al. (2024) discussed deep learning methods for event cameras [40], with just a brief subsection dedicated to 3D reconstruction and limited historical context. Therefore, an explicit, comprehensive, and updated survey is essential to systematically explore and evaluate the advancements in event-driven 3D reconstruction.
@@
As illustrated in Figure 1, Figure 3, and Figure 4, this paper systematically presents a hierarchical taxonomy and detailed chronological overview of event-based 3D reconstruction methods. In Section II, we introduce the fundamental working principles of event cameras, discussing (II-A) their differences from conventional cameras, (II-B) the event generation mechanisms, and (II-C) the typical event representations and output types commonly employed in reconstruction tasks. Sections III through V comprehensively review existing approaches categorised by camera setups and reconstruction methodologies. Specifically, Section III focuses on methods utilising stereo event cameras, divided into (III-A) geometry-based and (III-B) learning-based methods. Section IV examines monocular event-camera techniques, further subdivided into (IV-A) geometry-based methods producing traditional outputs, (IV-B) learning-based methods producing traditional outputs, (IV-C) Neural Radiance Fields-based methods, and (IV-D) 3D Gaussian Splatting-based methods. Section V addresses multimodal approaches integrating event cameras with complementary sensors, categorised into (V-A) methods with traditional outputs, (V-B) methods employing Neural Radiance Fields, and (V-C) methods using 3D Gaussian Splatting. Subsequently, Section VI compiles publicly available datasets relevant for benchmarking event-driven 3D reconstruction techniques. Finally, Section VII summarises this review, highlighting current research gaps and outlining promising future directions to advance the field.
II Event Camera with 3D Reconstruction
II-A Event Camera vs. Traditional Camera
Event cameras have several distinctive characteristics that set them apart from traditional frame-based cameras. They offer microsecond-level temporal resolution [11], enabling the capture of extremely fast motion without suffering from motion blur. With a dynamic range exceeding 120 dB [11], event cameras can handle both extremely bright and very dark scenes, making them highly suitable for environments with challenging or rapidly changing lighting conditions.
Referring to Figure 5 and Table I, unlike traditional cameras that capture full frames at fixed intervals, event cameras operate asynchronously by detecting changes in pixel intensity. Each pixel operates independently and triggers an event only when a significant change occurs [11], resulting in a continuous stream of spatio-temporal events. This principle leads to significantly lower data rates, reduced power consumption, and minimal latency, which are highly desirable characteristics for real-time applications and edge computing scenarios.
Figure 5: Event Camera vs. Traditional Camera
TABLE I: Compact Comparison between Event Cameras and Traditional Cameras
@@ II Event Camera with 3D Reconstruction
Time Resolution 1–10 $\mu$s/event 16–33 ms/frame (30–60 FPS)
Dynamic Range $>$120 dB 50–70 dB
@@
Latency $<$1 ms (end-to-end) 10–100 ms (sensor + processing)
@@
These unique characteristics make event cameras particularly well-suited for a variety of advanced computer vision tasks. Applications include, but are not limited to, high-speed object tracking [41], low-latency corner detection [42], real-time object recognition [43], and accurate depth estimation [44]. Event cameras have also shown great promise in video generation tasks [45, 46], light field video enhancement [47], and 3D reconstruction, where the combination of low latency and high temporal resolution contributes to precise and efficient scene understanding. With growing interest in neuromorphic vision and efficient visual sensing, event cameras are expected to play a key role in the future of robotics [48], AR/VR [49], autonomous driving [50], and beyond [26].
II-B Event Triggering & Event Representation
A mathematical approach is very common to explain the triggering of events. When the event camera detects a brightness change at a pixel $k$, it generates event data containing event coordinates $\mathbf{x}_{k}=(x_{k},y_{k})$, timestamp $t_{k}$ , and the polarity $p_{k}$. The brightness $L(\mathbf{x}_{k},t)=\log(I(\mathbf{x}_{k},t))$ is set as the pixel’s log intensity. The brightness change threshold $C$ usually varies by 10-15% [51]. An event $e_{k}=(\mathbf{x}_{k},t_{k},p_{k})$ is triggered when the brightness change $\Delta L$ at a pixel $k$ exceeds $C$, which can be expressed as:
@@ II-B Event Triggering & Event Representation
The event stream generated by an event camera is a continuous sequence of asynchronous and sparse data, where each event contains a timestamp, spatial coordinates, and polarity. This type of data is fundamentally different from conventional image frames. Due to its sparse and asynchronous nature, event data is difficult to directly extract features from and cannot be used as input to traditional deep learning models such as CNNs. Therefore, specialised preprocessing methods are typically required to convert event data into a more structured form, and this process is referred to as event representation. In addition to making the data more suitable for feature extraction, event representations help reduce data redundancy, improve computational efficiency, and enhance overall performance. Each representation method has its own strengths and limitations, and their applicability may vary depending on the task. Common event representation methods include the Event Frame, Time Surface, and Voxel Grid [52, 53, 54]. Event accumulate frame (EAF) and event gradient accumulation (EGA) are also very commonly used in current tasks with neural rendering output types [55, 56]. Previous surveys [26, 40] have already provided comprehensive overviews of these representations, and thus we do not elaborate further here.
II-C Event-based 3D Reconstruction Types
There are many tasks in 3D vision, but those considered as complete 3D reconstruction tasks typically require the final output to faithfully recover the spatial geometry of a scene or object and to be represented in an explicit and well-defined 3D format [57]. Such tasks go beyond estimating depth, disparity, or motion, which focus on reconstructing and modelling the full 3D structure of the environment. Specifically, the output should be a point cloud, mesh, voxel representation, an implicit 3D representation (such as SDF [58], NeRF [59], or Occupancy Networks [60]), etc., which can be directly used in downstream applications like rendering, path planning, robotic navigation, or virtual reality.
@@ II-C Event-based 3D Reconstruction Types
  • Point Cloud: The point cloud is one of the most straightforward 3D representations. A point cloud consists of a set of 3D coordinates $(x,y,z)$ [3], with each point potentially associated with additional attributes such as colour, normals, or timestamps [61, 62]. Point clouds can be obtained by back-projecting disparities estimated via stereo matching or depth estimation.
  • Voxel Model: Voxels discretise 3D space into regular cubic cells, where each voxel stores occupancy information or a probability value [5]. This representation is well-suited for volumetric modelling and is often used as input or supervision for neural networks.
  • Mesh Model: A mesh consists of vertices, edges, and faces that define the surface of a 3D object [4]. It is typically generated from point clouds or voxel representations through surface reconstruction techniques. Although the sparse nature of event data presents challenges for mesh generation, recent methods have demonstrated that surface reconstruction from event streams is feasible.
  • Depth Map: The depth map is a 2D image where each pixel encodes the depth value of the corresponding point in the scene [2]. Since a depth map is a 2D representation, it is less likely to be considered a complete 3D representation compared to the former three. However, with known camera intrinsics, a depth map can be easily converted into a point cloud, enabling more advanced surface modelling. Moreover, real-time depth estimation with event cameras is a topic closely related to, but distinct from, 3D reconstruction. Therefore, this survey includes some methods with 3D depth map output, and the following sections may cover crucial depth estimation works that produce 3D depth maps.
@@
While traditional outputs are suitable for explicit geometry modelling in tasks such as SLAM [63] and robotics [64], they often fall short in producing photorealistic renderings or modelling complex lighting and material properties. To address this, recent advances in neural rendering have introduced new 3D representations that are capable of modelling scenes in a continuous and differentiable manner [65]. These representations have enabled breakthroughs in novel view synthesis and high-fidelity scene reconstruction:
  • Neural Radiance Fields: Neural Radiance Fields (NeRF) is a neural network-based method for representing 3D scenes, proposed by Mildenhall et al. in 2020 [59]. It learns a continuous and differentiable 3D radiance field from multi-view 2D images and synthesises novel views [66], modelling the complex geometries and lighting effects [67, 68]. Since 2023, NeRF-based methods have been adapted for event-based 3D reconstruction [35]. A 3D scene can be reconstructed by applying differentiable volume rendering and extracting surfaces via iso-surface techniques. A NeRF models a continuous volumetric scene by training a multilayer perceptron (MLP) to learn a mapping from a 3D spatial location $\mathbf{x}\in\mathbf{R}^{3}$ and viewing direction $\mathbf{d}$ to the corresponding colour and volume density:
@@
  • 3D Gaussian Splatting (3DGS): 3D Gaussian Splatting, proposed by Kerbl et al. [69], represents a volumetric primitive in 3D space with attributes such as position, shape, orientation, and colour. In computer graphics, it serves as an explicit 3D representation that enables efficient differentiable rendering [70]. It provides an efficient solution that lies between traditional explicit point clouds and implicit volumetric rendering.
@@
    $f_{i}(\mathbf{p})=\sigma(\alpha_{i})\exp\left(-\frac{1}{2}(\mathbf{p}-\mu_{i})^{T}\Sigma_{i}^{-1}(\mathbf{p}-\mu_{i})\right)$ (5)
@@
III Methods with Stereo Event Cameras
Figure 7: Overview of methods with stereo event cameras. (a) Schematic of disparity matching. (b) Refocused DSI Fusion to replace disparity matching module, from Ghosh et al. [71]. (c) Self-Organising Maps from Steffen et al. [72]. (d) UNet-ResNet (CNN) from Nam et al. [73]
Stereo event cameras typically refer to two or more rigidly mounted event cameras. A stereo vision system composed of stereo event cameras is capable of asynchronously capturing event streams from the left and right viewpoints of a scene, enabling event-driven stereo matching and 3D scene reconstruction [74]. In general, the goal of stereo matching is to identify the most probable pairs of left and right events in both space and time, in order to compute disparity and subsequently estimate the depth structure of the scene.
Event-driven 3D-related tasks, such as stereo vision, were initially pioneered using stereo event cameras [75], and there is more research solving stereo vision [76, 31]. However, these works do not achieve a fully modelled 3D reconstruction as the final output, and therefore are not included in the survey.
Based on the data processing pipeline, 3D reconstruction methods using only stereo event cameras can be categorised into: (III-A) geometry-based methods and (III-B) learning-based methods. Table II and Figure 7 provide an overview of these methods. Figure 8 provides a visualisation of some of these methods producing depth map output.

### v3

@@
[…]
$2\qquad$ Rural Health Research Institute, Charles Sturt University, Orange, NSW, Australia. E-mail: Zhicheng Lu, zlu@csu.edu.au
Manuscript submitted: 2025-06; accepted: 2025-12
1 Introduction
Figure 1: Roadmap of 3D reconstruction with event cameras. It shows the development from event-based geometry to neural 3D rendering. With advances in technology, event-camera-based 3D reconstruction methods are achieving progressively higher accuracy and realism, enabling more complete and faithful 3D scene rendering.
@@ 1 Introduction
3D reconstruction is a fundamental technique that transforms 2D observations of real-world scenes, objects, or simulated environments into accurate 3D models [3dReconstruction]. Typical outputs from these reconstruction processes include depth maps [saxena2008make3d], point clouds [PointCloud], meshes [Mesh], and voxel representations [Voxel]. Numerous sensing technologies are used to capture the essential data required for 3D reconstruction, including conventional RGB cameras [RGB3drecon], RGB-D cameras [li2022high], structured light sensors [StructureLightSurvey], and LiDAR systems [LiDarSurvey, 3DSurvey]. However, each of these traditional sensors has inherent limitations: conventional RGB cameras struggle under extreme lighting conditions and rapid motion scenarios, whereas active sensors such as LiDAR and structured-light systems are typically bulky and require substantial physical space [3DSurvey]. Consequently, event cameras have emerged as an attractive alternative or complementary sensing modality, addressing many of these drawbacks.
Event cameras, also referred to as neuromorphic cameras, silicon retinas, or dynamic vision sensors, are bio-inspired sensors designed to asynchronously capture brightness changes rather than producing images at a fixed frame rate [lichtsteiner2008128]. Unlike conventional RGB cameras, each pixel in an event camera independently detects and reports intensity changes, effectively acting as an individual sensor element [posch2014retinomorphic]. When the brightness at a pixel exceeds a predefined threshold, the event camera generates event data consisting of the pixel’s coordinates, timestamp, and the polarity of the brightness change. Due to their unique sensing principle, event cameras exhibit outstanding advantages, such as high temporal resolution, low latency, and robustness to motion blur and challenging lighting conditions [lichtsteiner2008128]. These properties have driven extensive research on event cameras across various vision tasks, including object detection [Detection], recognition [Recognition], segmentation [Segmentation], and tracking [Tracking]; video enhancement tasks such as super-resolution [SuperResolution], frame interpolation [Interpolation], and deblurring [Deblurring]; as well as complex spatiotemporal modeling tasks including simultaneous localisation and mapping (SLAM) [SLAM], visual odometry (VO) [VO], and notably, 3D reconstruction. Event-based vision technologies have demonstrated considerable potential in diverse applications, ranging from autonomous driving [AutoDriving], robotics [Robot], and unmanned aerial vehicles (UAVs) [UAV] to industrial monitoring [Industry], etc.
@@
Leveraging their distinct sensing capabilities, event cameras have gained significant attention in recent years for performing 3D reconstruction. It is a domain traditionally dominated by standard cameras since the 1990s [[R1]]. Event-based 3D reconstruction approaches now span from occupancy mapping [[Towards]] and real-time dynamic scene reconstruction [[M3]], to achieving high-fidelity rendering of complex scenes [[G1]]. Figure 1 illustrates the roadmap and evolution of event-based 3D reconstruction techniques. Initial efforts utilising event cameras for 3D reconstruction began emerging prominently in the 2010s [[2013], [S100]]. Existing reconstruction strategies employ stereo and monocular event camera setups [[S9], [M1]], alongside multimodal approaches integrating event data with other sensory modalities [[DMT1]]. Recent advances further integrate sophisticated neural rendering methods, such as Neural Radiance Fields (NeRF) and 3D Gaussian Splatting (3DGS), to enable unprecedented levels of realism [[N3], [G1]].
Figure 3: Publication timeline in categories (from 2015 to May 2025). The research focus has shifted from early geometric and traditional learning methods to NeRF and 3D Gaussian Splatting, particularly in monocular and multimodal settings, where these emerging techniques have become dominant since 2023.
@@
Notably, a complete pipeline from raw event data acquisition to final 3D model reconstruction should be able to output a representable 3D structure, either in the form of a directly modellable 3D representation or at least a depth map that can be readily converted into a 3D point cloud (as in some depth estimation methods [[S4], [S0]]). However, numerous earlier works that lack a final 3D output, such as those in SLAM [SLAM], VO [VO], and 3D perception [[Non3d]], can be considered precursor steps toward 3D reconstruction, but should not be regarded as complete event-to-3D model reconstruction pipelines. In Figure 2, publications dedicated to event-based 3D reconstruction have significantly increased since 2015, experiencing rapid growth particularly in the last three years, underscoring the burgeoning interest and research value in this area. Since the choice of sensing configuration heavily influences the subsequent reconstruction pipeline, existing methods are typically categorised according to the camera setup: stereo, monocular, and multimodal approaches.
@@
Despite the steadily growing number of research publications, there currently exists no dedicated survey for event-driven 3D reconstruction. Three comprehensive but broader surveys have briefly touched upon aspects of 3D reconstruction using event cameras. Gallego et al. (2020) provided a wide-ranging overview of event cameras [[R1]], which included a brief section on 3D reconstruction. However, rapid advances have rendered parts of that review outdated. More recently, Chakravarthi et al. (2024) categorised event camera research tasks [[R2]], offering only limited insights into 3D reconstruction. Similarly, Zheng et al. (2024) discussed deep learning methods for event cameras [[R3]], with just a brief subsection dedicated to 3D reconstruction and limited historical context. Therefore, an explicit, comprehensive, and updated survey is essential to systematically explore and evaluate the advancements in event-driven 3D reconstruction.
Figure 4: Categorisation of methods in the survey. This survey distinguishes stero, monocular, and multimodal event camera systems by their inputs, and further categorises methods by processing type and output type into geometric approaches, learning-based methods, and neural rendering frameworks based on NeRF and 3D Gaussian Splatting.
@@
As illustrated in Figure 1, Figure 3, and Figure 4, this paper systematically presents a hierarchical taxonomy and detailed chronological overview of event-based 3D reconstruction methods. In Section 2, we introduce the fundamental working principles of event cameras, discussing (2.1) their differences from conventional cameras, (2.2) the event generation mechanisms, and (2.3) the typical event representations and output types commonly employed in reconstruction tasks. Sections 3 through 5 comprehensively review existing approaches categorised by camera setups and reconstruction methodologies. Specifically, Section 3 focuses on methods utilising stereo event cameras, divided into (3.1) geometry-based and (3.2) learning-based methods. Section 4 examines monocular event-camera techniques, further subdivided into (4.1) geometry-based methods producing traditional outputs, (4.2) learning-based methods producing traditional outputs, (4.3) Neural Radiance Fields-based methods, and (4.4) 3D Gaussian Splatting-based methods. Section 5 addresses multimodal approaches integrating event cameras with complementary sensors, categorised into (5.1) methods with traditional outputs, (5.2) methods employing Neural Radiance Fields, and (5.3) methods using 3D Gaussian Splatting. Subsequently, Section 6 compiles publicly available datasets relevant for benchmarking event-driven 3D reconstruction techniques, and Section 7 provides an overview of the most important and commonly used evaluation metrics in this field, categorized by type. Finally, Section 8 summarises this review, highlighting current research gaps and outlining promising future directions to advance the field.
2 Event Camera with 3D Reconstruction
2.1 Event Camera vs. Traditional Camera
Event cameras have several distinctive characteristics that set them apart from traditional frame-based cameras. They offer microsecond-level temporal resolution [lichtsteiner2008128], enabling the capture of extremely fast motion without suffering from motion blur. With a dynamic range exceeding 120 dB [lichtsteiner2008128], event cameras can handle both extremely bright and very dark scenes, making them highly suitable for environments with challenging or rapidly changing lighting conditions.
Referring to Figure 5 and Table 1, unlike traditional cameras that capture full frames at fixed intervals, event cameras operate asynchronously by detecting changes in pixel intensity. Each pixel operates independently and triggers an event only when a significant change occurs [lichtsteiner2008128], resulting in a continuous stream of spatio-temporal events. This principle leads to significantly lower data rates, reduced power consumption, and minimal latency, which are highly desirable characteristics for real-time applications and edge computing scenarios.
Figure 5: Event Camera vs. Traditional Camera. Traditional cameras output images at a fixed frame rate, whereas event cameras respond asynchronously to brightness changes in the scene, continuously generating a stream of events carrying spatial, temporal, and polarity information.
Table 1: Compact Comparison between Event Cameras and Traditional Cameras
@@ 2 Event Camera with 3D Reconstruction
[…]
Dynamic Range >>120 dB 50–70 dB
@@
Latency <<1 ms (end-to-end) 10–100 ms (sensor + processing)
@@
These unique characteristics make event cameras particularly well-suited for a variety of advanced computer vision tasks. Applications include, but are not limited to, high-speed object tracking [[35]], low-latency corner detection [[JDJC]], real-time object recognition [[36]], and accurate depth estimation [[37]]. Event cameras have also shown great promise in video generation tasks [[SPSC], qu2025evanimate], light field video enhancement [[LightField]], and 3D reconstruction, where the combination of low latency and high temporal resolution contributes to precise and efficient scene understanding. With growing interest in neuromorphic vision and efficient visual sensing, event cameras are expected to play a key role in the future of robotics [mahlknecht2022exploring], AR/VR [dong2024sevar], autonomous driving [maqueda2018event], and beyond [[R1]].
2.2 Event Triggering & Event Representation
A mathematical approach is very common to explain the triggering of events. When the event camera detects a brightness change at a pixel $k$, it generates event data containing event coordinates $\mathbf{x}_{k}=(x_{k},y_{k})$, timestamp $t_{k}$ , and the polarity $p_{k}$. The brightness $L(\mathbf{x}_{k},t)=\log(I(\mathbf{x}_{k},t))$ is set as the pixel’s log intensity. The brightness change threshold $C$ usually varies by 10-15% [[EMVS]]. An event $e_{k}=(\mathbf{x}_{k},t_{k},p_{k})$ is triggered when the brightness change $\Delta L$ at a pixel $k$ exceeds $C$, which can be expressed as:
@@ 2.2 Event Triggering & Event Representation
The event stream generated by an event camera is a continuous sequence of asynchronous and sparse data, where each event contains a timestamp, spatial coordinates, and polarity. This type of data is fundamentally different from conventional image frames. Due to its sparse and asynchronous nature, event data is difficult to directly extract features from and cannot be used as input to traditional deep learning models such as CNNs. Therefore, specialised preprocessing methods are typically required to convert event data into a more structured form, and this process is referred to as event representation. In addition to making the data more suitable for feature extraction, event representations help reduce data redundancy, improve computational efficiency, and enhance overall performance. Each representation method has its own strengths and limitations, and their applicability may vary depending on the task. Common event representation methods include the Event Frame, Time Surface, and Voxel Grid [TimeSurface, EventFrame, VoxelGrid]. Event accumulate frame (EAF) and event gradient accumulation (EGA) are also very commonly used in current tasks with neural rendering output types [[G3], [G13]]. Previous surveys [[R1], [R3]] have already provided comprehensive overviews of these representations, and thus we do not elaborate further here.
2.3 Event-based 3D Reconstruction Types
There are many tasks in 3D vision, but those considered as complete 3D reconstruction tasks typically require the final output to faithfully recover the spatial geometry of a scene or object and to be represented in an explicit and well-defined 3D format [geiger2011stereoscan]. Such tasks go beyond estimating depth, disparity, or motion, which focus on reconstructing and modelling the full 3D structure of the environment. Specifically, the output should be a point cloud, mesh, voxel representation, an implicit 3D representation (such as SDF [[SDF]], NeRF [[Nerf]], or Occupancy Networks [[occupancy]]), etc., which can be directly used in downstream applications like rendering, path planning, robotic navigation, or virtual reality.
@@ 2.3 Event-based 3D Reconstruction Types
  • Point Cloud: The point cloud is one of the most straightforward 3D representations. A point cloud consists of a set of 3D coordinates $(x,y,z)$ [PointCloud], with each point potentially associated with additional attributes such as colour, normals, or timestamps [deng2024multi, guo2021pct]. Point clouds can be obtained by back-projecting disparities estimated via stereo matching or depth estimation.
  • Voxel Model: Voxels discretise 3D space into regular cubic cells, where each voxel stores occupancy information or a probability value [Voxel]. This representation is well-suited for volumetric modelling and is often used as input or supervision for neural networks.
  • Mesh Model: A mesh consists of vertices, edges, and faces that define the surface of a 3D object [Mesh]. It is typically generated from point clouds or voxel representations through surface reconstruction techniques. Although the sparse nature of event data presents challenges for mesh generation, recent methods have demonstrated that surface reconstruction from event streams is feasible.
  • Depth Map: The depth map is a 2D image where each pixel encodes the depth value of the corresponding point in the scene [saxena2008make3d]. Since a depth map is a 2D representation, it is less likely to be considered a complete 3D representation compared to the former three. However, with known camera intrinsics, a depth map can be easily converted into a point cloud, enabling more advanced surface modelling. Moreover, real-time depth estimation with event cameras is a topic closely related to, but distinct from, 3D reconstruction. Therefore, this survey includes some methods with 3D depth map output, and the following sections may cover crucial depth estimation works that produce 3D depth maps.
@@
While traditional outputs are suitable for explicit geometry modelling in tasks such as SLAM [laidlow2019deepfusion] and robotics [banerjee2018robotic], they often fall short in producing photorealistic renderings or modelling complex lighting and material properties. To address this, recent advances in neural rendering have introduced new 3D representations that are capable of modelling scenes in a continuous and differentiable manner [yang2023jnerf]. These representations have enabled breakthroughs in novel view synthesis and high-fidelity scene reconstruction:
Figure 7: Overview of methods with stereo event cameras. (a) Schematic of disparity matching. (b) Refocused DSI Fusion to replace disparity matching module, from Ghosh et al. (c) Self-Organising Maps from Steffen et al. (d) UNet-ResNet (CNN) from Nam et al.
  • Neural Radiance Fields: Neural Radiance Fields (NeRF) is a neural network-based method for representing 3D scenes, proposed by Mildenhall et al. in 2020 [[Nerf]]. It learns a continuous and differentiable 3D radiance field from multi-view 2D images and synthesises novel views [qu2024nerf], modelling the complex geometries and lighting effects [zhuang2025simple, liu2025sem]. Since 2023, NeRF-based methods have been adapted for event-based 3D reconstruction [[N3]]. A 3D scene can be reconstructed by applying differentiable volume rendering and extracting surfaces via iso-surface techniques. A NeRF models a continuous volumetric scene by training a multilayer perceptron (MLP) to learn a mapping from a 3D spatial location $\mathbf{x}\in\mathbf{R}^{3}$ and viewing direction $\mathbf{d}$ to the corresponding colour and volume density:
@@
  • 3D Gaussian Splatting (3DGS): 3D Gaussian Splatting, proposed by Kerbl et al. [[Gaussian]], represents a volumetric primitive in 3D space with attributes such as position, shape, orientation, and colour. In computer graphics, it serves as an explicit 3D representation that enables efficient differentiable rendering [advance3DGS]. It provides an efficient solution that lies between traditional explicit point clouds and implicit volumetric rendering.
@@
    $\hskip-7.11317ptf_{i}(\mathbf{p})=\sigma(\alpha_{i})\exp\left(-\frac{1}{2}(\mathbf{p}-\mu_{i})^{T}\Sigma_{i}^{-1}(\mathbf{p}-\mu_{i})\right)$ (5)
@@
3 Methods with Stereo Event Cameras
Stereo event cameras typically refer to two or more rigidly mounted event cameras. A stereo vision system composed of stereo event cameras is capable of asynchronously capturing event streams from the left and right viewpoints of a scene, enabling event-driven stereo matching and 3D scene reconstruction [zhu2018realtime]. In general, the goal of stereo matching is to identify the most probable pairs of left and right events in both space and time, in order to compute disparity and subsequently estimate the depth structure of the scene.
Event-driven 3D-related tasks, such as stereo vision, were initially pioneered using stereo event cameras [mahowald1994vlsi], and there is more research solving stereo vision [[S3], [S100]]. However, these works do not achieve a fully modelled 3D reconstruction as the final output, and therefore are not included in the survey.
Based on the data processing pipeline, 3D reconstruction methods using only stereo event cameras can be categorised into: (3.1) geometry-based methods and (3.2) learning-based methods. Table 2 and Figure 7 provide an overview of these methods. Figure 8 provides a visualisation of some of these methods producing depth map output.